### v2

# The Influence of Code Smells in Efferent Neighbors on Class Stability

Abstract.
Understanding what drives code instability is essential for effective software maintenance, as unstable classes require larger or more frequent edits and increase the risk of unintended side effects. Although code smells are widely believed to harm maintainability, most prior stability studies examine only the smells within the class being modified. In practice, however, classes can change because their efferent neighbors (i.e., the classes they depend on) are modified due to ripple effects that propagate along static dependencies, even if the class itself is clean. Such ripple effects may be more severe when the efferent neighbor exhibits code smells. In addition, code smells rarely occur alone. They often appear together within a class or across classes connected by static dependencies, a phenomenon known as code smell interrelation. Such interrelation can lead to code smell interaction, where smells are directly connected through static dependencies and may further compound maintainability issues. However, the effect of code smell interrelation and interaction on code quality remains largely underexplored. Therefore, this study investigates whether the presence of code smells in a class’s efferent neighbors affects its stability, considering the factor of code smell interrelation and interaction. To achieve this, we mine one year of commit history from 100 top-starred GitHub projects, detect code smells and static dependencies, determine code smell interrelation and interaction, and model these factors as predictors of class stability.
1. Introduction
Changing source code is the heartbeat of software maintenance. For most maintenance tasks, such as fixing bugs, refactoring, or delivering new features, teams aim to keep edits small and localized because smaller changes cost less, disturb less context, and reduce the risk of unintended side effects (undef, ). A system exhibits low stability when parts of it require large edits during maintenance tasks or are frequently modified across multiple tasks (undefa, ; undefb, ). Such instability greatly increases the effort and risk of making changes. It is therefore important to understand which code characteristics drive this instability, enabling teams to address these factors, maintain system stability, and deliver changes effectively and efficiently.
One long-suspected factor is the presence of code smells: surface-level characteristics in source code that may indicate underlying design problems (undefc, ). Code smells are widely believed to degrade maintainability (undefc, ; undefd, ; undefe, ). Prior studies report that smelly classes tend to be edited more frequently and with greater change size than their non-smelly peers because of their large, complex, and non-cohesive nature (undefb, ; undeff, ; undefg, ; undefa, ). However, when developers change a class, it is not always because the class itself requires modification. A class may change because the classes it depends on via outgoing dependencies have changed, which can create ripple effects along static dependencies, even if the class itself is not smelly (undefh, ). We refer to these dependent classes as the class’s efferent neighbors [^1] Here, efferent is used as an adjective meaning “conducting outward” (i.e., outgoing), to emphasize that these neighbors are reached via outgoing dependencies from the class. Definition source: https://www.merriam-webster.com/dictionary/efferent. Figure 1 illustrates the concept: class $C$ has outgoing dependencies on classes $N_{1}$ and $N_{2}$, which we treat as $C$’s efferent neighbors. Changes in $N_{1}$ or $N_{2}$ can propagate to $C$ through these dependencies, creating ripple effects that may require $C$ to change as well. Consequently, when the efferent neighbors are smelly, these ripple effects may occur more frequently, thereby reducing the class’s stability. Focusing only on smells within the class captures the first situation but overlooks dependency-driven changes. To assess whether code smells also influence the stability of a class through this second pathway, we must analyze the smells present in its efferent neighbors and explicitly account for the static dependencies between classes.
Figure 1. Example of efferent neighbors: class $C$ depends on $N_{1}$ and $N_{2}$ via outgoing dependencies.
In addition, prior work suggests that code smells rarely occur in isolation. Multiple smells frequently appear together within the same class or in classes connected by static dependencies (undefi, ; undefj, ). This phenomenon is referred to as code smell interrelation, and multiple code smells can collectively exacerbate maintainability issues (undefk, ; undefl, ; undefj, ). A further consequence of this interrelation is code smell interaction, in which two smells are directly connected through static dependencies. These interactions may further complicate maintainability by affecting modularity, comprehensibility, and modifiability because of the direct structural connection between smelly artifacts (undefl, ; undefm, ). However, to the best of our knowledge, few studies have empirically examined how code smell interrelation and interaction influence code quality. Therefore, the goal of this confirmatory study is to assess the association between the presence of code smells in a class’s efferent neighbors and its stability, while accounting for the effects introduced by the presence of multiple code smells, that is, code smell interrelation and code smell interaction.
For this purpose, we will mine one year of commit history from 100 top-starred open-source Java projects. We will detect ten code smells, including five at the class level and five at the method level, and extract static dependencies to identify code smell interrelation and interaction. Each commit will be analyzed to obtain class-level stability measurements. We will model these data using negative binomial generalized linear models with project-level random intercepts and appropriate controls. Our expected contributions are as follows: (i) The design and execution of a large-scale empirical study, with all source code and analysis scripts to be made publicly available to enable full replication; and (ii) The first rigorous evidence on the influence of code smells in efferent neighbors on class stability, providing empirically grounded insights that can inform code smell detection, prioritization, and refactoring practices, and guide subsequent research on code smells and static dependencies.
2. Background and Related Work
2.1. Code Smells
Code smells are surface-level characteristics in source code that can indicate deeper design issues (undefc, ). There are different types of code smells, and they can be classified into class-level smells and method-level smells, as some affect the quality of a single method while others impact an entire class (undefn, ). In this study, we focus on ten types of code smells (undefn, ). The method-level smells are Feature Envy (FE), Brain Method (BM), Dispersed Coupling (DiCo), Intensive Coupling (IC), and Shotgun Surgery (SS). The class-level smells are God Class (GC), Brain Class (BC), Data Class (DC), Refused Bequest (RB), and Tradition Breaker (TB). Table 1 details their definitions (undefn, ).
Table 1. Definitions of Code Smells Considered in This Study (undefn, )
Code smell Definition
Feature Envy (FE) A method that uses another class’s data more than the data of its own class.
Brain Method (BM) A method that is excessively large and complex, uses many variables, and is deeply nested, tending to centralize too much functionality within a single class.
Dispersed Coupling (DiCo) A method that interacts with an excessive number of different classes by invoking only one or a few methods from each.
Intensive Coupling (IC) A method that invokes an excessive number of methods from one or a few other classes.
Shotgun Surgery (SS) A method that is invoked by too many other methods from many different classes.
God Class (GC) A large and complex class that concentrates much of the system’s functionality, frequently uses data from other classes, and has low cohesion, suggesting multiple responsibilities.
Brain Class (BC) A large and complex class that accumulates too much of the system’s functionality and contains at least one Brain Method. It exhibits low cohesion, but can be more cohesive than a God Class.
Data Class (DC) A low-complexity class that mainly stores data and provides access through public fields or accessor methods rather than encapsulating behavior.
Refused Bequest (RB) A subclass that makes little use of inherited protected members and rarely overrides parent methods.
Tradition Breaker (TB) A subclass that adds many new public methods while rarely specializing inherited ones, making its interface largely unrelated to its superclass.
Although artifacts may exhibit additional code smells beyond the ten we consider, our selection follows established practice in software engineering research. We focus on a subset of smells that (i) offer balanced coverage across both levels, (ii) include smells that are widely studied in software maintenance research, occur frequently in codebases, and have been linked to reduced maintainability (undefo, ; undefp, ; undefq, ; undefg, ; undefr, ; undefs, ; undeft, ), (iii) have clear definitions and widely accepted detection techniques (undefn, ), (iv) are supported by validated detection tools (undefu, ), and (v) align with prior studies, enabling meaningful comparison (undefb, ). Our goal is not to exhaustively cover all possible smells, but to analyze a representative subset. Consequently, we believe this choice makes our results more interpretable and practical.
We do not include a few well-known code smells such as Long Method (i.e., a method that is too long (undefc, )) and Large Class (i.e., a class that is too big (undefc, )) in this study. Although widely discussed, these smells are usually detected using a single aspect, such as method or class length, which on its own is often an overly sensitive heuristic (e.g., a Large Class that is fully cohesive and has a clear single responsibility, or a Long Method that is not deeply nested, accesses only a limited number of variables, and is easy to understand). In contrast, the smells we analyze are defined based on multiple aspects (e.g., the definitions of Brain Method and God Class include length and complexity, but also consider other aspects such as cohesion and nesting level), using richer structural and coupling-related characteristics that are more directly aligned with our dependency-based theory. Therefore, we focus on the ten well-defined smells to strengthen construct validity by aligning smell definitions with our dependency-based theory, while reducing detection noise from single-aspect, overly sensitive heuristics.
2.2. The Influence of Code Smells on Code Stability
Software stability is a sub-attribute of maintainability (undefv, ). Prior research has examined stability from different perspectives, each focusing on a particular category of software artifact. Salama et al. (undefh, ) classify these perspectives into four levels, where a level refers to the category of artifact being analysed. The four levels are the code level, the architecture level, the design level, and the requirements level. In this paper, we focus on the code level, as this is where code smells manifest; specifically, we are interested in the stability of class artifacts. At this level, stability is defined as (undefh, ): (i) intrinsic stability: the ability to remain largely unchanged over time (this captures an intrinsic property of the artifact itself, independent of external pressures such as ripple effects); and (ii) ripple-effect stability: the ability to resist the ripple effect of changes (an example of low ripple-effect stability is when one class changes and the classes depending on it also require modification).
Different metrics exist to measure code stability. Martin (undefw, ; undefx, ) proposes a package instability metric, which is proportional to a package’s efferent coupling, that is, the number of classes outside the package on which the classes inside the package depend. The rationale is that a package whose classes rely on many external classes is more likely to change when any of those dependent classes change. This conceptualization can be extended to the class level using the same rationale, where the efferent coupling of a class corresponds to the number of other classes it depends on (i.e., its efferent neighbors). In software evolution research, code stability is typically measured using two process metrics (undefa, ; undeff, ; undefb, ): change frequency, which measures how often a software entity is modified within a specific period, and change size, which quantifies the magnitude of each change and is usually expressed in terms of lines of code added or deleted (code churn) during each commit.
Empirical studies generally report that code smells are associated with lower code stability. Several investigations show that smelly classes tend to change more frequently and often with larger modifications, with this effect becoming stronger when multiple smells occur in the same class (undefb, ; undeff, ; undefg, ; undefa, ). Similar patterns were found in server-side web applications, where smelly files were more frequently updated and exhibited greater change size (undefy, ). A notable exception is reported by Olbrich et al. (undefd, ), who found that although God and Brain Classes exhibit higher change frequency and more defects in absolute terms, these differences disappear or reverse once class size is controlled for, indicating that size is a confounding variable in observed stability differences.
2.3. Architectural Smells and Their Evolution
Architectural smells are architecture-level design decisions or structural patterns that negatively impact system quality attributes (undefz, ). The key difference between architectural smells and code smells lies in the artifacts in which they manifest. Code smells are defined in terms of code-level (i.e., implementation-level) artifacts, such as methods and classes, whereas architectural smells are defined in terms of architectural-level artifacts, such as components and subsystems (undefz, ). For example, the code smell God Class and the architectural smell God Component (undefaa, ) both describe an artifact that has too many responsibilities. However, God Class refers to an overloaded class, while God Component refers to an overloaded component.
Previous research has studied the evolution of architectural smells. Sas et al. (undefab, ) tracked how individual smell instances evolve over time by mining three smell types across 14 open-source Java projects. They showed that smell types differ in multiple aspects, such as growth rate and lifetime, offering guidance for smell refactoring and prioritization. Extending this work, Sas et al. (undefac, ) analyzed how smells evolve in industrial systems. They tracked smells in nine C/C++ projects and complemented mining with interviews with 12 developers. Their findings suggest that smells can persist across releases, overlap within the same components, and be perceived by developers as contributing to maintainability difficulties. Developers also described how smell analysis can provide actionable refactoring insights.
Gnoyke et al. (undefad, ) studied the evolution of smells and their impact on system degradation by analyzing 14 open-source Java systems. They reported that smells remain mostly stable relative to code size, although some smell types have a greater impact on system degradation. They also identified properties that influence the strength of smells’ impact on system degradation. Recently, Gnoyke et al. (undefae, ) expanded this work by characterizing typical evolution patterns of smells. They found that some smells tend to accumulate and merge into more complex structures over time, while others remain relatively stable.
In this study, we focus on code smells because our goal is to analyze the stability of classes, which are code-level artifacts rather than architectural-level artifacts.
2.4. The Influence of Interrelated and Interacting Code Smells on Maintainability
Code smells are widely recognized and have been empirically shown to undermine maintainability (undefc, ; undefd, ; undefe, ). However, they rarely occur in isolation. Instead, multiple code smells often appear together in the same class or in related classes, collectively worsening maintainability. This phenomenon is known as code smell interrelation (undefl, ). Yamashita and Moonen (undefl, ) categorized interrelations as follows: (i) code smell collocation: multiple smells within the same class; and (ii) code smell coupling: smells appearing in different classes that are connected through static dependencies.
Several studies have examined how the collocation of multiple smells within the same class affects maintainability. Abbes et al. (undefk, ) found that individual smells showed little measurable impact on comprehension, whereas their collocation significantly increased the time, effort, and errors involved in understanding the code. Later work (undefaf, ; undefag, ) showed that classes with multiple collocated smells tend to have higher complexity and lower cohesion, with some combinations, such as a God Class collocated with Feature Envy, being especially detrimental. Developers also reported that collocated smells make code harder to understand and modify (undefaf, ).
Interestingly, prior work suggests that maintenance problems may arise not only from the presence of multiple interrelated smells, but from their interaction through static dependencies (undefah, ; undefj, ). In an exploratory study, Yamashita and Moonen (undefl, ) observed professional developers performing maintenance tasks and collected their reflections through interviews and think-aloud sessions. Developers indicated that collocated and coupled smells can interact along static dependencies, making code harder to follow and causing ripple effects during changes, which increase effort and defect risk. To the best of our knowledge, this remains the only study that directly links such interactions to maintainability, while later work has mainly identified recurring patterns of collocation and coupling without assessing their impact (undefj, ; undefai, ; undefaj, ; undefak, ; undefal, ).
Recently, Zhang et al. (undefm, ) formally defined the term code smell interaction: a code smell instance interacts with another code smell instance if there is a static dependency that directly connects the two instances. They concluded that such interactions are associated with degraded modularity.
The core difference between code smell coupling and code smell interaction is that coupling concerns dependencies between classes, while interaction concerns dependencies between individual code smell instances. Code smell instances can be more fine-grained than classes because some smells appear at the method level. For example, in Figure 2, class C1 contains method M1 and method CS1, where CS1 exhibits a method-level smell. Class CS2 contains method M2, where CS2 exhibits a class-level smell. The two classes are connected by a method-call dependency in both case A and case B, so code smell coupling exists for CS1 and CS2 in both cases. However, code smell interaction exists only in case B, because the call dependency directly connects method CS1 with class CS2.
Figure 2. Examples of interrelations and interactions between a method-level code smell instance CS1 (expressed as a blue square inside class C1) and a class-level code smell instance CS2 (expressed as the entire class colored in blue): (A) coupled but non-interacting code smells; (B) coupled and interacting code smells.
In summary, while prior studies have extensively examined the influence of collocated code smells on maintainability, few have addressed coupled code smells or the broader factor of interaction. This gap is particularly relevant to code stability. Focusing solely on how code smells within a class affect its stability addresses only intrinsic stability, while overlooking ripple-effect stability. When efferent neighbors contain code smells, they become unstable due to their large and non-cohesive nature. As a result, ripple effects along static dependencies may be triggered more frequently, requiring the assessed class to change more often as well. To investigate this, our study incorporates code smells in efferent neighbors and accounts for both code smell interrelation and interaction.
3. Study Design and Execution Plan
In this section, we outline the key components of our study. We begin by explaining the types of dependencies we consider, as they are fundamental for code smell coupling and interaction, followed by our research questions and hypotheses. We then present the variable definitions and our plan for obtaining and analyzing them.
3.1. Dependency
We utilize the following definition of dependency from Stevens et al. (undefam, ): “…A dependency is when the functioning of one element A requires the presence of another element B…” We interpret this as a directional connection between two code artifacts, flowing from the source artifact to the target artifact. We capture the requirement between the two through different relations. For example, a method (source artifact type) may create (relation type) a class (target artifact type). We therefore define a dependency type as a triple:
(relation type,source artifact type,target artifact type).($\text${relation type},\ $\text${source artifact type},\ $\text${target artifact type}).
In this paper, we consider five artifact types: class, interface, field, constructor, and method. Following and extending previous work (undefj, ; undefm, ), we focus on ten types of relations: call, create, contain, cast, use, throws, return, parameter, extend, and implement.
We treat all considered dependency types uniformly when deriving dependency-based measurements and conducting analyses. Specifically, dependencies of different types contribute equally; we do not apply dependency-type-specific weights. We further discuss the implications of this decision in Section 4.
3.2. Research Questions and Hypotheses
In this section, we list our research questions (RQ) and formalize the hypotheses for each RQ. For all hypotheses, the null hypothesis ($H_{0}$) is that there is no statistically significant effect or association between the independent variable and the stability measure. The alternative hypothesis ($H_{a}$) is that the specified relationship is significant. Below, we list only the alternative hypotheses.
Additionally, stability will be assessed based on change frequency and change size (see Section 3.4). Due to space constraints, each hypothesis is worded in terms of stability in general, but in the study, each one will be evaluated as two separate hypotheses.
We use the term focal class to refer to the class whose stability we are measuring, and the term efferent neighbors to refer to all the classes it depends on. When we test the effect of code smell coupling and interaction (RQ3 and RQ4), we examine both (a) the effect without accounting for the presence or number of code smells in the focal class and its efferent neighbors, and (b) the effect with these factors held constant. We do this because (i) deconfounding is necessary, since code smells in the focal class or its neighbors may independently decrease stability, and (ii) we want to determine whether coupling or interaction adds additional instability beyond what can be explained by the smells already present in those locations. This allows us to distinguish the raw association from the unique contribution of coupling or interaction after controlling for baseline smell conditions.
3.2.1. RQ1
Although the influence of collocated code smells on stability has been studied extensively, we include this aspect for completeness and to enable comparison with previous literature. Thus, our RQ1 and its hypotheses are:
RQ1: Does the existence of code smells in a class influence its stability?
$\boldsymbol{H1.1_{a}}$: Smelly classes are significantly less stable than non-smelly classes.
$\boldsymbol{H1.2_{a}}$: There is a significant negative association between the number of code smells a class exhibits and its stability.
$\boldsymbol{H1.3_{a}}$: There is a significant negative association between the variety of code smells a class exhibits and its stability.
3.2.2. RQ2
As Martin (undefw, ; undefx, ) suggests, instability is proportional to efferent coupling. Our theory is that if the efferent neighbors exhibit code smells, the focal class becomes less stable than if the efferent neighbors are clean. This is because prior empirical studies suggest that smelly classes tend to change more frequently (undefb, ; undeff, ; undefg, ; undefa, ), which in turn may lead the classes depending on them to change more often. Such a situation might even be true for non-smelly classes. Therefore, we test both the population of all classes and the population of non-smelly classes to isolate the effect of smells in the focal class. Hence, our RQ2 and its hypotheses are:
RQ2: Does the existence of code smells in a class’s efferent neighbors influence its stability?
$\boldsymbol{H2.1_{a}}$: Classes with smelly efferent neighbors are significantly less stable than classes without.
$\boldsymbol{H2.2_{a}}$: There is a significant negative association between the number of code smells in a class’s efferent neighbors and its stability.
$\boldsymbol{H2.3_{a}}$: There is a significant negative association between the variety of code smells in a class’s efferent neighbors and its stability.
$\boldsymbol{H2.4_{a}}$: Non-smelly classes with smelly efferent neighbors are significantly less stable than those without.
$\boldsymbol{H2.5_{a}}$: There is a significant negative association between the number of code smells in a non-smelly class’s efferent neighbors and its stability.
$\boldsymbol{H2.6_{a}}$: There is a significant negative association between the variety of code smells in a non-smelly class’s efferent neighbors and its stability.
3.2.3. RQ3
If both the focal class and at least one of its efferent neighbors exhibit code smells, we refer to this as efferent code smell coupling  [^2] Here, efferent modifies code smell coupling as defined in Section 2.4, indicating that there exists an outgoing dependency that leads to code smell coupling. We speculate that this situation may be worse for stability than cases where (a) only the focal class is smelly, (b) only the efferent neighbors are smelly, or (c) neither is smelly. Our theory is that when a code smell instance in an efferent neighbor changes, the ripple effect through the dependency may also require developers to change the code smell instance in the focal class. The potential need to change code smell instances in both the focal class and an efferent neighbor at once may lead to more frequent and larger changes to the focal class, thereby further reducing its stability. Hence, our RQ3 and its hypotheses are:
RQ3: Does the existence of efferent code smell coupling influence a class’s stability?
$\boldsymbol{H3.1_{a}}$: Classes with efferent code smell coupling are significantly less stable than classes without, if the presence of code smells in the class itself and in its efferent neighbors are ignored.
$\boldsymbol{H3.2_{a}}$: Classes with efferent code smell coupling are significantly less stable than classes without, if the presence of code smells in the class itself and in its efferent neighbors are held constant.
3.2.4. RQ4
We employ the term efferent code smell interaction  [^3] Here, efferent modifies code smell interaction as defined in Section 2.4, indicating that there exists an outgoing dependency that leads to code smell interaction to refer to a situation in which a code smell instance CS1 interacts with another code smell instance CS2, where CS1 resides in the focal class, CS2 resides in its efferent neighbor, and there exists at least one dependency in which the CS1 instance is the source artifact and the CS2 instance is the target artifact. We refer to such a dependency as an efferent code smell interaction dependency. For a focal class, we define the number of efferent code smell interactions as the count of all such code smell pairs, and we define efferent code smell interaction intensity as the number of efferent code smell interaction dependencies.
We speculate that this situation may further influence stability. This case arises when the focal class and an efferent neighbor both contain code smells (i.e., efferent code smell coupling), and the two smell instances are directly connected by a static dependency where the focal class is the source artifact. In such cases, when a smell in the efferent neighbor changes, the dependency makes it more likely that developers must also change the corresponding smell in the focal class. As a result, developers are more likely to need to modify smell instances in two classes at once. In contrast, in efferent code smell coupling without a direct static dependency between the smell instances, the probability of needing to change smells in both classes at once is lower. Additionally, since interacting code smells are associated with more static dependencies between them (undefm, ), there are more routes for changes to ripple into the focal class, making it more likely to be changed more often and more extensively. Hence, our RQ4 and its hypotheses are:
RQ4: Does the existence of efferent code smell interaction influence a class’s stability?
$\boldsymbol{H4.1_{a}}$: Classes with efferent code smell interactions are significantly less stable than classes without, when the number of code smells they exhibit and the number of code smells in their efferent neighbors are ignored.
$\boldsymbol{H4.2_{a}}$: There is a significant negative association between the number of efferent code smell interactions in a class and its stability, when the number of code smells it exhibits and those in its efferent neighbors are ignored.
$\boldsymbol{H4.3_{a}}$: There is a significant negative association between the intensity of efferent code smell interactions in a class and its stability, when the number of code smells it exhibits and those in its efferent neighbors are ignored.
$\boldsymbol{H4.4_{a}}$: Classes with efferent code smell interactions are significantly less stable than classes without, when the number of code smells they exhibit and the number of code smells in their efferent neighbors are held constant.
$\boldsymbol{H4.5_{a}}$: There is a significant negative association between the number of efferent code smell interactions in a class and its stability, when the number of code smells it exhibits and those in its efferent neighbors are held constant.
$\boldsymbol{H4.6_{a}}$: There is a significant negative association between the intensity of efferent code smell interactions in a class and its stability, when the number of code smells it exhibits and those in its efferent neighbors are held constant.
3.3. Independent Variables (IVs)
We define the following independent variables (IVs):
  • Is the focal class smelly (IsSmelly): Indicates whether the focal class exhibits at least one code smell.
  • Number of code smells in the focal class (#SmellFoc): The number of code smell instances in the focal class.
  • Variety of code smells in the focal class (VarSmellFoc): The number of different types of code smells present in the focal class.
  • Has smelly efferent neighbor (HasSmellEff): Indicates whether at least one efferent neighbor of the focal class exhibits a code smell.
  • Number of smells in efferent neighbors (#SmellEff): The total number of code smell instances across all efferent neighbors of the focal class.
  • Variety of smells in efferent neighbors (VarSmellEff): The number of different types of code smells present across all efferent neighbors of the focal class.
  • Has efferent code smell coupling (HasEffCoup): Indicates whether the focal class exhibits efferent code smell coupling.
  • Has efferent code smell interaction (HasEffInt): Indicates weather the focal class exhibits efferent code smell interaction.
  • Number of efferent code smell interactions (#EffSmellInt): The number of code smell pairs that exhibit efferent code smell interaction for the focal class.
  • Efferent code smell interaction intensity (EffIntInten): The number of efferent code smell interaction dependencies.
Since our unit of analysis is the class, we attribute both class-level and method-level smell instances to their enclosing class (i.e., each method-level smell instance is assigned to the class that contains the method). We treat smell instances and smell types uniformly (i.e., no weighting by smell level or smell type); each detected smell instance is aggregated in the analysis and contributes equally to our smell-based IVs.
The rationale for measuring the number and variety of code smells is that higher values are often associated with degraded maintainability, particularly reduced stability (undefg, ; undefb, ). The rationale for measuring interaction intensity is that higher intensity reflects a greater number of dependencies connecting smelly artifacts, which can harm comprehensibility and compound a developer’s ability to effectively perform modifications (undefl, ).
3.4. Dependent Variables (DVs)
In this section, we introduce the metrics used to measure class stability, which serve as our dependent variables (DVs). For each focal class, we measure its stability over the observation period using the following metrics:
  • Change Frequency (ChF): The number of commits that modify the class during the period.
  • Change Size (ChS): The total number of lines of code added or deleted in the class, excluding blank lines and comments, aggregated over all commits during the period.
These two metrics capture complementary aspects of stability: ChF reflects how often a class changes, while ChS reflects how much it changes. Using both provides a more complete view of class stability.
Since stability is measured with two metrics, each hypothesis listed in Section 3.2 will be treated as two separate hypotheses. We add an additional superscript to indicate which metric the hypothesis is testing. For example, $\boldsymbol{H1.1_{a}}$ will be treated as two hypotheses: $\boldsymbol{H1.1_{a}^{\text{{ChF}}}}$ and $\boldsymbol{H1.1_{a}^{\text{{ChS}}}}$.
3.5. Control Variables (CVs)
For all hypotheses, we include the following control variables (CVs):
  • Class size (ClSize): The lines of code in the class, excluding comments and blank lines.
  • Number of efferent neighbors (#EffNei): The number of classes on which the focal class depends (i.e., its efferent coupling).
The rationale for including these controls is to reduce confounding. Larger classes inherently have a higher probability of being changed (undefd, ), and classes with more efferent neighbors are also more likely to change because modifications to any of their efferent neighbors may trigger changes in the focal class (undefw, ; undefx, ).
For hypothesis $\boldsymbol{H3.2}$, we additionally control for IsSmelly and HasSmellEff. For hypotheses $\boldsymbol{H4.4}$ to $\boldsymbol{H4.6}$, we additionally control for #SmellFoc and #SmellEff. The rationale is explained in Section 3.2. We do not include VarSmellFoc and VarSmellEff as control variables in hypotheses $\boldsymbol{H4.4}$ to $\boldsymbol{H4.6}$. These variables capture smell-type diversity and are likely to be correlated with #SmellFoc and #SmellEff. Including both count and variety may introduce multicollinearity and reduce interpretability. Because the primary purpose of the additional controls is to adjust for baseline smell conditions, we treat #SmellFoc and #SmellEff as reasonable proxies for the main deconfounding goal.
Table 2 lists the IVs and CVs for each hypothesis. The DVs are always the two listed in Section 3.4.
Table 2. Independent and Control Variables per Hypothesis
Hypotheses Independent Variables Control Variables
$\boldsymbol{H1.1}$ IsSmelly ClSize, #EffNei
$\boldsymbol{H1.2}$ #SmellFoc ClSize, #EffNei
$\boldsymbol{H1.3}$ VarSmellFoc ClSize, #EffNei
$\boldsymbol{H2.1}$ & $\boldsymbol{H2.4}$ HasSmellEff ClSize, #EffNei
$\boldsymbol{H2.2}$ & $\boldsymbol{H2.5}$ #SmellEff ClSize, #EffNei
$\boldsymbol{H2.3}$ & $\boldsymbol{H2.6}$ VarSmellEff ClSize, #EffNei
$\boldsymbol{H3.1}$ HasEffCoup ClSize, #EffNei
$\boldsymbol{H3.2}$ HasEffCoup ClSize, #EffNei, IsSmelly, HasSmellEff
$\boldsymbol{H4.1}$ HasEffInt ClSize, #EffNei
$\boldsymbol{H4.2}$ #EffSmellInt ClSize, #EffNei
$\boldsymbol{H4.3}$ EffIntInten ClSize, #EffNei
$\boldsymbol{H4.4}$ HasEffInt ClSize, #EffNei, #SmellFoc, #SmellEff
$\boldsymbol{H4.5}$ #EffSmellInt ClSize, #EffNei, #SmellFoc, #SmellEff
$\boldsymbol{H4.6}$ EffIntInten ClSize, #EffNei, #SmellFoc, #SmellEff
3.6. Repository Mining
For each system under consideration, we first select a snapshot that has at least one year of subsequent commit history, which serves as the observation window. For each class in that version, we obtain its size and the number of efferent neighbors, which serve as the CVs (ClSize and #EffNei). We then detect code smells in the class and its efferent neighbors, and obtain the measurements for code smell coupling and interaction. These constitute the IVs. Next, for each commit (including direct commits and pull-request merges) in the one-year period on the main branch, we identify all classes modified by the commit and compute the amount of code change for each class. Specifically, we obtain the diff for the class and count the added and deleted lines of code after removing blank lines and comments. Aggregating these values across all commits yields the two DVs (ChF and ChS) for the class.
We assume a one-to-one mapping while mining the commit history: (i) the focal class in the snapshot can be matched to exactly one class throughout the mining period, and (ii) each commit’s changes can be attributed to that same single focal class in the snapshot over time. This assumption ensures that the IV and CVs measured for a focal class at the snapshot can be unambiguously associated with its DV computed from aggregated changes, and that each DV value can be unambiguously associated with the IV and CVs of a single focal class in the snapshot.
Classes that are renamed or moved do not violate this assumption; they are treated as the same evolving class and are retained in the analysis. We exclude classes involved in class merge or class split refactorings during the observation period because these operations violate the one-to-one mapping, either by making the focal class no longer unambiguously attributable to the entity whose changes are aggregated (i.e., class split), or by making the entity whose changes are aggregated no longer unambiguously attributable to a single focal class in the snapshot (i.e., class merge). We do not apply special handling to other refactoring commits; such commits do not violate our assumption and are intentionally included because our stability measures are designed to capture all changes, regardless of whether they arise from feature work, bug fixes, or refactoring.
The rationale for choosing a one-year time span is to avoid large changes in the function of the class between the time of code smell detection and the commit-mining period, since such changes may alter the class’s intrinsic stability.
It is possible that new smells are introduced or that smells are refactored from the focal class or its efferent neighbors during the period, which may influence code smell coupling and interaction. Our goal is to investigate class instability influenced by smells, smell couplings, and smell interactions, rather than explaining why smells were introduced or removed. Therefore, we do not assume that smells, smell coupling, or smell interaction remain constant during the period.
3.7. System Selection
We will conduct our study on top-starred open-source projects from GitHub. The rationale for using top-starred projects is to prioritize the most popular and widely adopted projects (undefb, ). By ranking projects in descending order of the number of stars, we will select the first 100 projects that meet the following inclusion criteria (IC), following recommendations from Kalliamvakou et al. (undefan, ): IC1. More than 100 forks (indicates meaningful developer interest). IC2. At least 20 contributors (to avoid personal or low-activity projects). IC3. At least one year of commit history with at least 50 commits within that one-year period (to ensure meaningful development activity, approximately one commit per week). IC4. More than 80% of the code written in Java (due to the smell detection tool limitation). IC5. Not an educational project (educational projects may not represent real software development processes).
To further contextualize the selected systems and to support the adequacy of the one-year observation window, we will report descriptive statistics on maintenance activity within the one-year period. Specifically, for each system, we will compute the number of commits in the selected one-year window and system-level churn (e.g., lines added/deleted aggregated over the window), and summarize these quantities across all selected systems using the minimum, maximum, median, mean, and standard deviation. Although our inclusion criteria aim to ensure meaningful development activity, these post-hoc statistics provide a transparent empirical characterization of the actual activity observed in the analysis window.
3.8. Data Analysis
In this section, we describe how we analyze the data. The analysis methods and statistical models discussed in this section apply to all our RQs and hypotheses.
It is reported that most software engineering metrics do not follow a normal distribution (undefao, ). Therefore, we assume our IVs, DVs, and CVs are non-normal. We use statistical methods that do not assume normality and apply log transformations when appropriate. Due to the large-scale nature of our dataset, normality tests become overly sensitive; therefore, even though we assume non-normality, we verify practical deviations by examining quantile-quantile plots.
We analyze our DVs using generalized linear models (GLMs). GLMs extend ordinary linear regression to non-normal outcomes by using an appropriate distribution and a link function to relate predictors to the expected outcome (undefap, ). In empirical software engineering, regression-based models such as GLMs are widely used to model a variety of software engineering data (undefaq, ; undefar, ). In our analysis, for each hypothesis, we fit a separate GLM with the corresponding DV (either ChF or ChS) as the response and the IV and CVs specified for that hypothesis in Table 2 as predictors.
Both DVs are count data (i.e., non-negative integers arising from counting) and are modeled using negative binomial GLMs with a log link. Poisson and negative binomial GLMs are both standard for count data, but Poisson assumes equidispersion (the mean equals the variance) (undefap, ). Because previous work suggests software engineering count measures can be over-dispersed relative to this assumption, we adopt the negative binomial model for its flexibility (undefas, ). For completeness, we also fit Poisson GLMs and compute the dispersion statistic to confirm the presence of overdispersion.
Since we assume non-normality, we apply log transformations to our count-based IVs (ClSize, #EffNei, #SmellFoc, #SmellEff, VarSmellFoc, VarSmellEff) before modeling, to reduce the influence of extreme values and obtain smoother relationships (undefat, ). We do not transform the DVs, as they are modeled directly using appropriate GLMs.
A random effect is a model component that captures group-level variability, and it is increasingly used in empirical software engineering (undefau, ; undefav, ; undefaw, ). We include a project-level random intercept in all models to account for systematic differences between projects, such as variation in age, development activity, commit practices, and team size. This adjustment ensures that comparisons between classes are not biased by project-specific factors while allowing the estimated effects of code smells to generalize across projects (undefax, ). Using a random intercept rather than fixed effects is appropriate given the large number of projects and our aim to estimate population-level effects that generalize beyond the specific projects, rather than fitting a separate parameter for each project (undefau, ).
For each fitted model, we will report the regression coefficient ($\beta$) for each IV together with its standard error, 95% confidence interval, and $p$-value (computed using Wald tests for the regression coefficients; one-sided as specified by each directional hypothesis), which collectively describe the estimated direction of the effect, its precision, and its statistical significance. To convey the practical magnitude of effects (i.e., effect size), we will report Incidence Rate Ratios (IRRs) (undefay, ) and Average Marginal Effects (AMEs) (undefaz, ). These translate the estimated coefficients into interpretable changes in complementary ways: IRRs present effects as percentage (multiplicative) changes, while AMEs present effects as average raw-unit (additive) changes. Model performance will be summarized using the log-likelihood (undefap, ) and McFadden’s pseudo-$R^{2}$ (undefaaa, ), both of which provide standard goodness-of-fit evaluations for GLMs. Together, these reporting elements ensure a comprehensive assessment of model adequacy.
We evaluate hypotheses using one-sided tests, consistent with the directional predictions specified in the study design. Because each RQ is addressed via multiple hypotheses, evaluated for both stability outcomes (ChF and ChS), and each RQ corresponds to a distinct confirmatory claim about a different theoretical mechanism linking code smells to stability (e.g., mechanisms related to smells in the class, in efferent neighbors, efferent code smell coupling, and efferent code smell interaction), we define the family of tests separately for each RQ. The study is not designed to evaluate a single claim that “code smells are related to stability” in general, but rather to test whether such a relationship is supported under specific mechanisms; accordingly, inferential statements are interpreted per RQ, and conclusions for one RQ are not used to support claims about another. We therefore control the false discovery rate using the Benjamini-Hochberg (BH) procedure separately within each RQ, applying BH to the set of one-sided p-values associated with the IV in the regression model for every hypothesis under that RQ, resulting in adjustments over 6, 12, 4, and 12 p-values for RQ1, RQ2, RQ3, and RQ4, respectively. Statistical significance is assessed at $\alpha=0.05$: an alternative hypothesis is considered accepted if the coefficient ($\beta$) of the IV is in the predicted direction and its one-sided BH-adjusted $p$-value ¡ 0.05; otherwise, we conclude there is insufficient evidence to accept it. Both unadjusted and BH-adjusted p-values will be made available.
3.9. Implementation
Projects will be filtered (according to the first four inclusion criteria discussed in Section 3.7), selected, and downloaded automatically using the GitHub API (undefaab, ). Inclusion Criterion 5 (i.e., not an educational project) will be enforced manually by inspecting each project repository, primarily the project documentation (e.g., the README). Code smells will be detected using JSpIRIT (undefaac, ; undefaad, ). The rationale for choosing this tool is that it can detect the ten code smells we will study, and it strictly follows the definitions and detection strategies provided by Lanza and Marinescu (undefn, ), ensuring transparency. Static dependencies, efferent neighbors of the focal class, and class size will be detected and measured using the static analysis tool CodeQL (undefaae, ). This information will be collectively analysed using Python to obtain all IVs and CVs. Git commands (undefaaf, ) will be used to analyse commit history and extract information such as diffs, and RefactoringMiner (undefaag, ) will be used to identify class refactorings, including class rename, move, split, and merge. Python will automate the process and compute the DVs for each included class. Finally, R will be used to fit the GLMs via the glmmTMB package (undefaah, ) and to obtain all statistical outputs.
4. Threats to Validity
In this section, we present what we believe are the most critical threats.
A threat to internal validity stems from the temporal mismatch between measuring code smells and dependencies at a single snapshot and measuring class stability over a subsequent one-year observation window. Because code smells and dependencies may evolve during the window, predictor values measured at the initial snapshot may differ from those at later points in time. We mitigate this threat by selecting an observation window that is long enough to capture meaningful maintenance activity, but not so extended that predictors are likely to evolve substantially. However, since we cannot fully rule out time-varying changes in predictors within the window, our results should be interpreted as associations between snapshot conditions and subsequent stability rather than evidence of causal effects.
Another threat to internal validity is our choice of a one-year commit history as the observation window, since different time spans may produce different stability patterns. We selected a one-year window rather than the full project lifecycle because our goal is to investigate how code smells in a class’s efferent neighbors relate to the class’s subsequent stability, rather than to characterize how class stability evolves across the entire project lifecycle. Additionally, as the time span grows longer, it becomes more likely that (1) historical behavior and age-related lifecycle effects are captured (i.e., as a project becomes older, its development and maintenance practices may evolve), (2) a class’s functionality changes substantially, and (3) the temporal mismatch discussed in the previous paragraph becomes more severe. These factors can confound the relationship between predictors and the measured stability outcomes. Shorter periods, in contrast, may not include enough maintenance activity to meaningfully assess stability. To mitigate the risk of insufficient maintenance activity, our system selection criteria require a certain level of development activity during the observation window, and we will report descriptive statistics (e.g., commits and churn) to empirically characterize the maintenance activity during the one-year period.
Another threat to internal validity is that the studied projects may differ systematically in many factors, such as age and commit practices. These factors may lead to different baseline levels of class stability independent of the studied predictors. To mitigate this risk, we include a project-level random intercept to account for project-specific baseline differences. This adjustment helps ensure that class-level comparisons are made on a more comparable basis across projects, and supports findings that generalize beyond any single project.
A further threat to internal validity concerns our project selection process. Following the recommendations of Kalliamvakou et al. (undefan, ), we exclude personal, inactive, and educational repositories and focus on widely adopted projects to maximize insights into real code evolution. However, we limit our analysis to Java systems due to the language support of the selected smell-detection tool. While this procedure reduces noise introduced by low-quality or non-representative repositories, it also limits the extent to which our results generalize to non-Java or less popular projects.
A threat to construct validity is the discrepancy between tool-detected code smells and the design problems perceived by developers. While subjective judgments can never be fully captured, we reduce this discrepancy by using a tool that faithfully implements the widely adopted smell definitions of Lanza and Marinescu (undefn, ). Because these definitions underpin many refactoring guidelines, smell-detection tools, and code-smell studies, they approximate a community-level consensus among experienced developers. As a result, the detector’s output is more likely to align with their expectations than tools that rely on opaque or non-standard detection rules. Nevertheless, individual developers may still disagree with specific detections depending on domain knowledge, design intent, or business logic, and our findings should be interpreted with this limitation in mind.
Another threat to construct validity is that we treat all dependency types uniformly (i.e., without dependency-type-specific weighting) when defining efferent neighbors and deriving interaction measures, although different dependency types may induce ripple effects with different strengths. This decision is motivated by the lack of prior evidence to justify dependency-type-specific weighting; applying such weighting would introduce subjective design choices. Consequently, our results should be interpreted as aggregate effects across dependency types, rather than dependency-type-specific mechanisms.
A further threat to construct validity is that we do not include Duplicated Code (i.e., multiple similar or identical code fragments that appear in different locations (undefaai, )) as a smell in this study. Our theory focuses on ripple effects that propagate along explicit static dependencies between classes, whereas changes related to duplicated code can spread through clone similarity and clone-management practices (e.g., copy-paste) without explicit dependency links. Consequently, clone-related propagation mechanisms may not be captured by our dependency-based measurements, and our findings should be interpreted as applying to dependency-driven rather than clone-driven change propagation.
A threat to conclusion validity concerns the suitability of our statistical analysis. To verify distributional assumptions, Poisson GLMs will be fitted and dispersion statistics will be computed; if the statistics indicate overdispersion, negative binomial GLMs will be used. Negative binomial GLMs are well-suited for overdispersed count data and will be implemented via the established glmmTMB package in R. Because model selection will be based on explicit distributional checks and will rely on well-validated methods, it is unlikely that the resulting conclusions will be affected by the choice of model or software.
5. Data Availability and Ethical Considerations
We will manage and share research data (including, but not limited to, our source code, scripts, and processing pipeline, as well as repository-derived data such as project source code, commit history, and analysis outputs produced by running the pipeline) in accordance with the University of Auckland Research Data Management Policy  [^4] https://www.auckland.ac.nz/en/about-us/about-the-university/policy-hub/research-innovation/research-data-management/research-data-management-policy.html [^5] https://research-hub.auckland.ac.nz/managing-research-data/ethics-integrity-and-compliance/research-data-management-policy-guidance. Accordingly, we will make the research data publicly available via an appropriate repository to support open science and replicability, unless it cannot be published due to legal, ethical, data sovereignty, or commercial constraints.
We acknowledge that repository data may contain person-related information (e.g., developer identifiers). We will ensure that data collection and processing comply with relevant ethical and legal requirements. Actions such as removing or anonymizing personally identifiable information will be taken to protect individual privacy, ensuring compliance with those requirements and enabling data sharing in support of open science.